MEDCON6 Shutdown Plan

From GlueXWiki
Jump to: navigation, search


Hall D is going into a state consistent with the expectation of no beam for the remainder of the year.

The plan is to have minimum people in the Hall or Counting House simultaneously. The 6' social distancing rule shall be adhered to. If not possible, AD (Rolf) approval is required and PPE is needed.

The shift worker/RC will start powering down electronics in conjunction with the detector experts after the last run is complete, at about 9 am.


If it is possible, all work done in the hall should be done after RADCON surveys the Hall, collimator cave and Tagger. Mark Stevens contacted Dave Hamlette and it looks like we will get RADCON support around noon Tuesday. If using rapid/Controlled access, only one person in the labyrinth at a time. The Run Coordinator in conjunction with the Work Coordinator will coordinate these efforts to minimize people in the Hall.

Once in restricted access, the Work Coordinator or Designee will direct traffic in the hall

General Timing of Work in the Hall

It is assumed the Hall will be in rapid (controlled) access until about noon

  • Mark Dalton, Mark Stevens and Tim Whitlatch will be in about 8:00 am Tuesday
  • Nick Sandoval comes in around 9:30 to shut things down in the hall
  • Lubomir comes in around 9:30 to take care of his responsibilities
  • Keith Blackburn and Chris Allen come in around 10am to start working Hall D proper beamline vacuum, then move on to collimator and Tagger when surveyed
  • Josh Foyles and Bobby Bunton come in around 11am to change Argon bank, then proceed to do chiller work
  • Chris Keith comes in 1pm to secure Target
  • Keith B and Chris A will proceed to DIRC work and continue the next morning if needed
  • Scot Spiegel will come in Wednesday morning 9am to secure power supplies and the gas system


  1. EMPTY the target. This boils almost all the hydrogen back into the storage tank and takes about 10 minutes or so; -- Shiftworker
  2. When the target reads "EMPTY and ready for beam", turn the pulse tube refrigerator OFF. -- Tim
  3. Text Chris Keith at 757-746-9277. -- Tim
    • Chris will come in around 1pm Tuesday and secure the system. That should require an access to the hall of only 5 minutes;
      • Ensure that both H2 tanks are valved off ?
      • Turn off the turbo station that is pumping on the OVC ?
      • Turn off the pulse tube compressor and kill power at the service disconnect ?


Intention is to keep magnet full and cold at 4k

  1. Set ramp goal to 0 A and ramp down -- Shift worker or Tim
  2. Cryogenics is secured for the shutdown -- Tim and Cryo
  3. Install a roughing pump on coil 4 chimney and turn on - keep roughing valve closed -- Keith/Chris
    1. this is in case cryo plant has issues and magnet is warmed to 80K -- Tim
    2. If magnet is warming to 80K, all 4 coil vacuum valves should be opened to the manifold -- Tim
  4. After magnet PS reaches 0A, turn off power supply and depress mushroom button on PS -- Scot
  5. Stop recording PXI data -- Tim


  1. Thin window guards
    1. Tagger window shutter lowered - Keith B/Chris A
    2. Reinstalled in Collimator Cave - Chris A/Keith B
    3. Reinstalled on Pair Spectrometer - Keith B/Chris A
    4. Reinstall on FCAL beampipe - Keith B/Chris A
  2. Beamline Vacuum (all valves closed)
    1. Goniometer pump staying on - Keith B/Chris A
    2. Tagger pump staying on - Keith B/Chris A
    3. Turbo cart for 10" beampipe between collimator on tagger remains On
    4. All Beam line valves closed (and disconnected)
    5. Collimator Cave beam line turn pumps off and bleed up per procedure - Keith B/Chris A
    6. Pair Spectrometer pump OFF - Keith B/Chris A
    7. FCAL platform beam pipe vacuum off and bled up - Keith B/Chris A
  3. Cooling
    1. Fan on microscope OFF - Josh/Bobby
    2. BCAL Chillers OFF (drain and flush with nitrogen)- Josh/Bobby
    3. BCAL readout nitrogen purge - reduce to 3 l/m -- Tim
    4. FDC Chiller OFF - Josh/Bobby
    5. Blower CDC OFF - Josh/Bobby
  4. Drift Chamber gas
    1. Lower gas flow
      • FDC -- Lubomir
      • CDC -- change to 1/2 flow - Scot
    2. Change over Argon bank - Josh/Bobby
    3. Bypass alcohol bubbler -- Scot
  5. Turn off Tagger NMR probe - Tim
  6. Pair spec and Sweep Magnets
    • Ramp down pair spec magnet, turn off power supply and depress mushroom button. - Scot
    • Ramp down sweep magnet, turn off power supply and depress mushroom button. - Scot
  7. DIRC
    1. Turn off water purification system -- Keith B/Chris A
    2. Drain both Optical boxes and holding tank (RadCon support) - Keith B/Chris A
      • Obtain water samples from North OB, South OB, and two from storage tank: leave samples in Hall for later analysis. -- Keith B/Chris A
      • RadCon approval required before emptying water to drain. Note: storage tank doesn't have enough capacity to drain both OB.
    3. Turn off blowers (only after LV supply is disconnected) -- Nick Sandoval
    4. Increase nitrogen purge to 100 l/hr for the OBs -- Keith B/Chris A
    5. Drain water skid and add nitrogen purge
    6. Continue nitrogen purge at normal rate for the Bar Boxes
  8. ComCal
    1. Turn off chiller and drain water - Josh/Bobby
    2. Keep nitrogen purge on detector at 15 l/hr
  9. Flood gates - do not install at this time
  10. TEDF Cleanroom - no work going on there, leave as is, lights out
  11. BDX Tent - once HV is turned off remotely, turn off crates and Bertha - Mark S.
  12. Record status of this checklist into Logbook (HDLOG)- Mark S


Retract devices, Computers ON, DAQ Crates OFF, control crates ON

All this work that can be done on the CSS screens should be performed by the Shift leader/Mark Dalton if not otherwise noted

  1. Secure for shutdown
    1. Radiators retracted
    2. Profile monitor retracted
    3. Collimator in blocked position
    4. TPOL converter retracted
    5. PS convertor retracted
    6. COMPCAL retracted
    7. TAC retracted
  2. Detector Power Status: OFF
    1. BCAL (LV, bias)
    2. CDC (HV, LV)
    3. DIRC (HV)
      • LV must definitely be off since cooling will be disabled -- Nick Sandoval
    4. FCAL. [Requires access to turn OFF: first hit the crash button near the door, then power off the connection boxes (on the back) and the power strip on the front.] -- Nick Sandoval
    5. CCAL
    6. FDC (HV, LV)-- Lubomir
    7. GEM/TRD (HV, LV) -- Lubomir
    8. PS, PSC
    9. TAGH -- Nick Sandoval
    10. TAGM -- Nick Sandoval
    11. TOF (HV, LV) Lubomir
    12. ST (Bias, LV) Lubomir
    13. Beam line Voltages (Active Target, TAC, halo)
    14. TPOL
  3. Keep Power ON
    1. All computers, switches, control modules
    2. Accelerator crates (US1-1, US1-2)
  4. Crate Power OFF
    1. DAQ crates: VXS, VME Hall D
    2. DAQ crates: VXS, VME Tagger building
  5. Record status in the Log book.


This is was initially based on the shutdown checklists.

File:HCO - Hall D Electronics Hall D Shutdown Checklist.pdf
File:HCO - Hall D Mechanical Hall D Shutdown Checklist.pdf
File:HCO - Hall D Solenoid Hall D Shutdown Checklist.pdf
File:HCO - ENP Target Hall D Shutdown Checklist.pdf